# Bayesian inferences on covariant density functionals from multimessenger astrophysical data: Nucleonic models

### v2

@@ 9 $\epsilon\,(\rho)$ MeV fm-3 N3LO G
In Eqs. (9), we truncate the expansion at different orders in density in the isoscalar and isovector sectors, while also presenting higher-order coefficients for each sector—specifically, $Z_{\rm sat}$, $K_{\rm sym}$ and $Q_{\rm sym}$. These coefficients are, in fact, predictions derived from the lower-order parameters, because the latter uniquely fix the parameters of our CDF. The correlation matrix among these coefficients, shown in Fig. 2, is obtained from the posterior distribution using only the symmetric nuclear matter (SNM) constraints at saturation density, as listed in the upper panel of Table 2. The results indicate that most of these coefficients exhibit little to no correlation, except for the pair $(M^{\ast}_{\rm D},\,Q_{\rm sat})$, which still shows a relatively weak correlation of approximately 0.3. Importantly, this framework allows for the independent variation of specific macroscopic characteristics within their acceptable ranges while keeping others fixed. This enables a systematic investigation of how these variations influence the EoS of dense matter and the properties of CSs [96, 97].
@@ III Bayesian inference framework
Bayesian analysis allows us to infer the probability distribution of unknown parameters in a model by exploiting the information in observed data. This is accomplished via the Bayes’ theorem,
@@
In our process of sampling, for each sample of parameters in Eq. (11) we first calculate the remaining five non-free parameters $(b_{\sigma},\,c_{\sigma},\,a_{\omega},\,b_{\omega},\,c_{\omega})$ in the isoscalar sector from Eq. (4) with a reference density $\rho_{\rm ref}$. Given these twelve parameters and $\rho_{\rm ref}$ we can readily compute the seven nuclear matter characteristics at saturation density listed in Table 2. The value of $\rho_{\rm ref}$ is updated each time, simultaneously with the five on-free parameters, in an iterative process, until the requirement $\rho_{\rm ref}=\rho_{\rm sat}$ is reached. At this stage, we have determined all the relevant parameters for the prior samples. We then compute the core EoS of stellar matter by incorporating the additional conditions of weak equilibrium and charge neutrality, which govern the composition of CSs. To ensure consistency, we smoothly match our core EoS to the crust EoS from Refs.[98, 99] at the crust-core transition density $0.5\,\rho_{\rm sat}$. Finally, the global properties of CSs, including mass, radius, and TD, are obtained by integrating the Tolman-Oppenheimer-Volkoff equations [100, 101] and the associated equations for TD [102, 103].
In the present analysis, we employ the Python implementation of the Bayesian inference tool MULTINEST [104, 105, 106], which makes use of the Nested Sampling technique, to generate the posterior samples and estimate the marginalized distributions. The distributions presented for each run are counted from posterior samples of $\sim 3\times 10^{4}$ EoSs to ensure the robustness of the inferred results. Below, we present a detailed discussion on the prior and likelihood we adopted in the current analysis.
@@ III.1 Nuclear matter properties
The coefficients of the lower-order terms in the expansions (9), specifically $E_{\rm{sat}}$, $K_{\rm{sat}}$ and $J_{\rm{sym}}$ , are well constrained. Additionally, the Dirac effective mass $M^{\ast}_{\rm D}$ at saturation density — crucial for an accurate description of finite nuclear phenomena, including spin-orbit splitting — is also well constrained. In contrast, the higher-order coefficients, such as $Q_{\rm{sat}}$ and $L_{\rm{sym}}$ remain less well determined, though they serve as asymptotic boundary conditions for the parameterization of CDF models. For parameters known to within approximately $10\%$, we assume Gaussian prior distributions, while for those with greater uncertainties, we adopt uniform priors. The mean values and standard deviations, or the respective parameter intervals, are summarized in Table 2. Notably, we consider broader ranges for $Q_{\text{sat}}$ and $L_{\text{sym}}$ to encompass the possible values suggested by various works [107, 108, 109, 110, 111, 81, 91].
The Gaussian likelihood function for parameters i is defined as
@@
where $d_{i}$ and $\sigma_{i}$ are the data and its standard uncertainties, respectively, and $\mu_{i}$ the corresponding model value. For quantities $Q_{\rm{sat}}$ and $L_{\rm{sym}}$, the likelihood functions are implemented as Heaviside step functions of the form
@@
In addition to the constraints above, we incorporate results for pure neutron matter based on chiral effective field theory ($\chi$EFT) interactions which constrain the low-density regime of nucleonic EoS. In practice, we utilize the N3LO calculations from Ref. [73] to extract the energy per particle and pressure at densities of 0.08, 0.12, and 0.16 fm-3, which serve as constraints in our analysis. We further assume that the associated uncertainties follow a Gaussian distribution at the 68.3% confidence level. This approach allows us to retain models that fall outside the strict $\chi$EFT band, rather than discarding them outright. It is worth noting that the uncertainty band reported in Ref. [73] lies within the broader range of results obtained from various $\chi$EFT interactions compiled in Ref. [84].
@@
Those CDF parameterizations that are compatible with the $\chi$EFT are soft at low densities and remain such up to $2$-$3\,\rho_{\rm sat}$. More quantitatively, they have the range of symmetry energy and its slope at nuclear saturation density within the range $29.5\leq J_{\rm sym}\leq 33.5$ MeV and $35\leq L_{\rm sym}\leq 65$ MeV (at 95.4% CI), respectively. This leads to small radii for canonical $1.4\,M_{\odot}$ mass stars, $R\lesssim 13.5$ km, as shown in Fig. 4. More details on the posteriors of nuclear matter parameters are discussed in Subsec. IV.2.
@@ III.2.1 NICER data
To date, the NICER collaborations have delivered the joint measurement of mass and radius through pulse profile modeling of four millisecond pulsar: a massive $\sim 2.1\,M_{\odot}$ star PSR J0740+6620 and two canonical-mass $\sim 1.4\,M_{\odot}$ stars PSR J0030+0451 and PSR J0437-4715, and a low-mass $\sim 1\,M_{\odot}$ one PSR J1231-1411. We construct our likelihood function for each of the sources using the Gaussian kernel density estimation (KDE) with the released posterior $(M,\,R)$ samples $\bm{S}_{i}$,
@@
where the mass $M$ and radius $R$ for the star are functions of its central pressure and of the sampled EoS parameters. For present analysis, we implement the following measured $(M,\,R)$ samples as shown in Fig. 1 and listed below. The final likelihood is the product of four likelihoods for each of these sources.
@@
  • PSR J0437-4715 (J0437). We use the first mass and radius estimates for this brightest pulsar by using the 2017-2021 NICER X-ray spectral-timing data from Ref. [43]. The preferred CST+PDT model used informative priors on mass, distance, and inclination from PPTA radio pulsar timing data and taking into account constraints on the non-source background and validated against XMM-Newton data [43].
@@
  • PSR J1231-1411 (J1231). The inference results obtained with the preferred PDT-U model show a strong sensitivity to the choice of radius priors, with stable and likely converged outcomes achieved only under constrained radius priors. We consider the mass and radius estimates from the preferred PDT-U model with limited radius priors given in Ref. [44]. One estimate limited the radius to be consistent with the previous observational constraints and EoS analyses (model (i)), and the other estimate used an uninformative prior for which the radius was limited to ranges 10-14 km (model (ii)).
@@ III.2.2 GW data
where $S(f)$, $d(f)$, and $h(\bm{\theta}_{\rm GW},f)$, respectively, represent the power spectral density of the detector noise, the detected strain signal, and the expected strain from a waveform model. The vector $\bm{\theta}_{\rm GW}$ includes parameters that are particularly useful to infer EoS properties, $\bm{\theta}^{\rm EoS}_{\rm GW}$, and nuisance parameters, $\bm{\theta}^{\rm nuis.}_{\rm GW}$, that are essential when performing analysis on GW-emitting binaries. Including a dozen of nuisance parameters, however, significantly slows down the sampling process and hence we calculate the likelihood through the high-precision interpolation in TOAST [113] that marginalizes over these parameters,
@@
where $\mathcal{M}=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ is the chirp mass, $q=M_{1}/M_{2}$ is the mass ratio of the event $i$, and $\Lambda_{1}(M_{1})$ and $\Lambda_{2}(M_{2})$ the TD of the individual star, respectively. The TDs are related to the component star’s mass and radius through the underlying EoS and its radial profile.
@@ III.2.4 HESS J1731-347
In the case of unusually light neutron star HESS J1731-347 (hereafter J1731), a consensus on its gross parameters has not been reach yet on the basis of observational modeling. Nevertheless, we tentatively incorporate this source into our analysis for theoretical exploration. Specifically, we adopt one of the two most reliable results from Ref. [64], which utilized the most advanced physical modeling and the most precise distance measurements available. Similar to the NICER sources, we apply the KDE method to the released model samples to generate the posterior distributions, which are then treated as the likelihood $\mathcal{L}_{\rm{HESS}}(\bm{\theta}_{\rm{EoS}})$.
Figure 5: Posteriors for mass-radius (upper panels) and mass-tidal deformability (lower panels) distributions under different astrophysical scenarios explored in this work. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B (left panels) and F (right panels). The various color solid lines show the results at 95.4% CI for scenarios which take additionally into account the mass-radius estimates of HESS J1731 (B2 and F2), the mass measurement of PSR J0952 (B3 an F3), or both of these masurements (B4 and F4). Finally, the dash-dotted lines correspond to the scenarios which do not include the $\chi$EFT constraints. In the lower panels, the insets magnify the TDs for typical masses of CSs involved in GW events. In the upper panels, the elliptical contours of the input mass-radius estimates of HESS J1731 and the band for the mass of PSR J0952 are shown as well for reference.
Figure 6: Probability distribution functions (PDFs) for maximum mass $M_{\rm max}$ under different astrophysical scenarios. The shaded regions represent the distributions for scenarios B1 (left panel) and F1 (right panel), while the various color solid lines correspond to the scenarios which take additionally into account the mass measurement of PSR J0952 (B2 an F2), mass-radius estimates of HESS J1731 (B3 and F3), or both of these two measurements (B4 and F4). The dash-dotted lines correspond to neglecting the $\chi$EFT constraints in scenarios B1 and F1.
Figure 7: Correlation between the radii $R_{2.0}$ and $R_{1.4}$ (upper panels), and the maximum mass $M_{\rm max}$ and the radius-difference $\Delta R=R_{1.4}-R_{2.0}$ (lower panels) under different astrophysical scenarios. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B1 (left panels) and F1 (right panels), while the various color solid lines correspond to 95.4% CI for the scenarios which take additionally into account the mass measurement of PSR J0952 (B2 an F2), mass-radius estimates of HESS J1731 (B3 and F3), or both of these two measurements (B4 and F4). The dash-dotted line for scenarios (B1 and F1) correspond to neglecting the $\chi$EFT constraints.
@@ IV Multimessenger inferences on compact stars and dense matter
We apply the theoretical and astrophysical constraints both selectively and in concert to elucidate the importance of the individual impact of the ultra-compact low-mass star (HESS J1731) and the very high mass star (PSR J0952) as well as the overall picture that emerges from a combination of such constraints.
Before embarking onto detailed discussion we present the scenarios that we consider in Table 4. Specifically, we will examine two baseline scenarios, labeled B1 and F1, which share the same data for PSR J0437 and PSR J0740 but incorporate two different analyses for PSR J0030 using different models of the surface temperature patterns, and two different analyses for PSR J1231 using different radius priors. For each baseline scenario, we implement further the data for the ultra-compact low-mass star HESS J1731 (denoted as scenarios B2 and F2), and for the high-mass star PSR J0952 (denoted as scenarios B3 and F3). Finally we take into account both the two constraints (denoted as scenarios B4 and F4) in our analysis.
@@ IV.1 Implications for properties of compact stars
Figure 8: Posteriors for the nucleonic compact-star EoS (upper panels) distribution and the proton fraction (lower panels) under different scenarios. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B1 (left panels) and F1 (right panels), while the various color solid lines correspond to 95.4% CI for the scenarios which take additionally into account the mass measurement of PSR J0952 (B2 and F2), the mass-radius estimates of HESS J1731 (B3 and F3), and both of these two measurements (B4 and F4). The dash-dotted line for scenarios (B1 and F1) correspond to neglecting the $\chi$EFT constraint. In each panel, the contours show the corresponding distributions of the respective $2.0\,M_{\odot}$ and the maximum-mass configurations for scenarios B1 (B4) and F1 (F4). In lower panels the orange band labeled $F_{\rm DU}$ shows the admissible threshold values for the onset of direct Urca (DU) cooling process.
@@
Incorporating the mass and radius estimates for HESS J1731 results in only a slight shift—approximately 0.1 km—to the lower-radius side in the posterior region for scenario B2 (F2) compared to B1 (F1); see Fig. 5. As expected, the posterior distributions for tidal deformability ($\Lambda$) exhibit similar trends to the $M$-$R$ posteriors due to the strong positive correlation between $\Lambda$ and $R$ for a given mass (e.g. Ref [47] and references therein). Consequently, the posterior region overlaps with the mass and radius estimates for HESS J1731 only at the 95.4% confidence level. Meanwhile, the impact on the maximum mass ($M_{\rm max}$) distribution remains negligible in scenarios B1 and B2 (F1 and F2); see Fig. 6. This suggests that the mass and radius measurements of four pulsars from NICER, combined with the TDs inferred from GW170817, GW190425 and the constraints from $\chi$EFT calculations, primarily define the posterior region of the $M$-$R$ distribution. The effect of the $\chi$EFT constraint on the $M$-$R$ posterior is particularly evident in scenarios B1 and F1 (Fig. 5), where it imposes a lower bound on the radii of sub-canonical-mass stars. Additionally, the mass and radius inferences derived for HESS J1731 shows only weak overlap with the $M$-$R$ ellipse for PSR J1231 (Fig. 1), which limits the influence of HESS J1731 data in distinguishing viable EoS.
@@
Including the mass measurement for PSR J0952 instead, namely for scenarios B3 (F3), the posterior region is shifted to a larger radius region by 0.1-0.4 km compared to scenario B1 (F1). A more significant difference appears for stars with mass $M>1.4\,M_{\odot}$. This is because the radii for sub-canonical-mass stars are well-constrained, while the NICER inferences for PSR J0030 and J0740, allow for radii larger than $14$ km, which then allow stiffer high-density EoS. As anticipated, the maximum mass $M_{\rm max}$ is enlarged by $0.1\,M_{\odot}$, with, in particular, a higher probability density for scenarios F3; see Fig. 6. The maximum mass now is $M_{\rm max}=2.28_{-0.19}^{+0.20}\,M_{\odot}$ and $2.39_{-0.23}^{+0.19}\,M_{\odot}$ respectively for scenarios B3 and F3. Meanwhile, the distributions of maximum mass $M_{\rm max}$ is somewhat narrowed compared to scenarios B1 and F1, respectively.
@@
Finally, incorporating both HESS J1731 and PSR J0952 data, namely for scenarios B4 and F4, the boundaries for the posterior regions are located well within the range set by the two cases discussed above. The distributions of maximum mass $M_{\rm max}$ are rather close to those of scenarios B3 and F3, respectively. To conclude, the $M$-$R$ estimates for HESS J1731 and mass measurement for PSR J0952 affect only the finer details of the posterior compared to the case where these constraints are not included.
@@
Reference [114] pointed out that the sign of the difference of the radii of CSs of masses 1.4 and $2.0\,M_{\odot}$, $\Delta R=R_{1.4}-R_{2.0}$ is an indicator that the EoS softens (if positive) or stiffens (if negative) at high densities. Figure 7 illustrates the posterior distributions for the correlation between $R_{1.4}$ and $R_{2.0}$, and the correlation between the maximum mass $M_{\rm max}$ and the radius difference $\Delta R$. It is seen that the correlations between these parameters are similar across all scenarios. For scenario B, the preference is for $\Delta R$ being positive, i.e., $R_{1.4}>R_{2.0}$, but for scenarios F, negative $\Delta R$ is allowed even at 68.4% CI. More details of correlations among properties of CSs will be discussed below in Subsec. IV.3. The predicted key quantities of CSs under different astrophysical scenarios are summarized in the Appendix.

### v3

@@ 9 $\epsilon\,(\rho)$ MeV fm-3 N3LO G
In Eqs. (9), we truncate the expansion at different orders in density in the isoscalar and isovector sectors, while also presenting higher-order coefficients for each sector—specifically, $Z_{\rm sat}$, $K_{\rm sym}$ and $Q_{\rm sym}$. These coefficients are, in fact, predictions derived from the lower-order parameters, because the latter uniquely fix the parameters of our CDF. The correlation matrix among these coefficients, shown in Fig. 2, is obtained from the posterior distribution using only the symmetric nuclear matter (SNM) constraints at saturation density, as listed in the upper panel of Table 2. The results indicate that most of these coefficients exhibit little to no correlation, except for the pair $(M^{\ast}_{\rm D},\,Q_{\rm sat})$, which still shows a relatively weak correlation of approximately 0.3 (Pearson coefficient) . Importantly, this framework allows for the independent variation of specific macroscopic characteristics within their acceptable ranges while keeping others fixed. This enables a systematic investigation of how these variations influence the EoS of dense matter and the properties of CSs [95, 96].
@@ III Bayesian inference framework
Bayesian analysis allows us to infer the probability distribution of unknown parameters in a model by exploiting the information in observed data. This is accomplished via Bayes’s theorem,
@@
In our process of sampling, for each sample of parameters in Eq. (11) we first calculate the remaining five non-free parameters $(b_{\sigma},\,c_{\sigma},\,a_{\omega},\,b_{\omega},\,c_{\omega})$ in the isoscalar sector from Eq. (4) with a reference density $\rho_{\rm ref}$. Given these twelve parameters and $\rho_{\rm ref}$ we can readily compute the seven nuclear matter characteristics at saturation density listed in Table 2. The value of $\rho_{\rm ref}$ is updated each time, simultaneously with the five non-free parameters, in an iterative process, until the requirement $\rho_{\rm ref}=\rho_{\rm sat}$ is reached. At this stage, we have determined all the relevant parameters for the prior samples. We then compute the core EoS of stellar matter by incorporating the additional conditions of weak equilibrium and charge neutrality, which govern the composition of CSs. To ensure consistency, we smoothly match our core EoS to the crust EoS from Refs. [97, 98] at the crust-core transition density $0.5\,\rho_{\rm sat}$. Finally, the global properties of CSs, including mass, radius, and TD, are obtained by integrating the Tolman-Oppenheimer-Volkoff equations [99, 100] and the associated equations for TD [101, 102].
In the present analysis, we employ the Python implementation of the Bayesian inference tool MULTINEST [103, 104, 105], which makes use of the Nested Sampling technique, to generate the posterior samples and estimate the marginalized distributions. The distributions presented for each run are counted from posterior samples of $\sim 3\times 10^{4}$ EoSs to ensure the robustness of the inferred results. Below, we present a detailed discussion of the prior and likelihood we adopted in the current analysis.
@@ III.1 Nuclear matter properties
The coefficients of the lower-order terms in the expansions (9), specifically $E_{\rm{sat}}$, $K_{\rm{sat}}$ and $J_{\rm{sym}}$ , are well constrained. Additionally, the Dirac effective mass $M^{\ast}_{\rm D}$ at saturation density — crucial for an accurate description of finite nuclear phenomena, including spin-orbit splitting — is also well constrained. In contrast, the higher-order coefficients, such as $Q_{\rm{sat}}$ and $L_{\rm{sym}}$ remain less well determined, though they serve as asymptotic boundary conditions for the parametrization of CDF models. For parameters known to within approximately $10\%$, we assume Gaussian prior distributions, while, for those with greater uncertainties, we adopt uniform priors. The mean values and standard deviations, or the respective parameter intervals, are summarized in Table 2. Notably, we consider broader ranges for $Q_{\text{sat}}$ and $L_{\text{sym}}$ to encompass the possible values suggested by various works [106, 107, 108, 109, 110, 81, 90].
The Gaussian likelihood function for parameter i is defined as
@@
where $d_{i}$ and $\sigma_{i}$ are the datum and its standard uncertainty, respectively, and $\mu_{i}$ the corresponding model value. For quantities $Q_{\rm{sat}}$ and $L_{\rm{sym}}$, the likelihood functions are implemented as Heaviside step functions of the form
@@
In addition to the constraints above, we incorporate results for pure neutron matter based on chiral effective field theory ($\chi$EFT) interactions which constrain the low-density regime of nucleonic EoS. In practice, we utilize the N3LO calculations from Ref. [73] to extract the energy per particle and pressure at densities of 0.08, 0.12, and 0.16 fm-3, which serve as constraints in our analysis. We further assume that the associated uncertainties follow a Gaussian distribution at the 68.3% CI. This approach allows us to retain models that fall outside the strict $\chi$EFT band, rather than discarding them outright. It is worth noting that the uncertainty band reported in Ref. [73] lies within the broader range of results obtained from various $\chi$EFT interactions compiled in Ref. [83].
@@
Those CDF parametrizations that are compatible with the $\chi$EFT are soft at low densities and remain so up to $2$-$3\,\rho_{\rm sat}$. More quantitatively, they have the valus of symmetry energy and its slope at nuclear saturation density within the range $29.5\leq J_{\rm sym}\leq 33.5$ MeV and $35\leq L_{\rm sym}\leq 65$ MeV (at 95.4% CI), respectively. This leads to small radii for canonical-mass $1.4\,M_{\odot}$ stars, $R_{1.4}\lesssim 13.5$ km, as shown in Fig. 4. More details on the posteriors of nuclear matter parameters are discussed in Subsec. IV.2.
@@ III.2.1 NICER data
To date, the NICER collaborations have delivered the joint measurement of mass and radius through pulse profile modeling of four millisecond pulsars: a massive $\sim 2.1\,M_{\odot}$ star PSR J0740+6620 and two canonical-mass $\sim 1.4\,M_{\odot}$ stars PSR J0030+0451 and PSR J0437-4715, and a low-mass $\sim 1\,M_{\odot}$ one PSR J1231-1411. We construct our likelihood functions for each of the sources using the Gaussian kernel density estimation (KDE) with the released posterior $(M,\,R)$ samples $\bm{S}_{i}$,
@@
where the mass $M$ and radius $R$ for the star are functions of its central pressure and of the sampled EoS parameters. For the present analysis, we implement the following measured $(M,\,R)$ samples as shown in Fig. 1 and listed below. The final likelihood is the product of four likelihoods for each of these sources.
@@
  • PSR J0437-4715 (J0437). We use the first mass and radius estimates for this brightest pulsar by using the 2017-2021 NICER X-ray spectral-timing data from Ref. [43]. The preferred CST+PDT model used informative priors on mass, distance, and inclination from PPTA radio pulsar timing data, taking into account constraints on the non-source background and validating against XMM-Newton data [43].
@@
  • PSR J1231-1411 (J1231). The inference results obtained with the preferred PDT-U model show a strong sensitivity to the choice of radius priors, with stable and likely converged outcomes achieved only under constrained radius priors. We consider the mass and radius estimates from the preferred PDT-U model with limited radius priors given in Ref. [44]. One estimate limited the radius to be consistent with the previous observational constraints and EoS analyses (model (i)), and the other estimate used an uninformative prior for which the radius was limited to range 10-14 km (model (ii)).
@@ III.2.2 GW data
where $S(f)$, $d(f)$, and $h(\bm{\theta}_{\rm GW},f)$, respectively, represent the power spectral density of the detector noise, the detected strain signal, and the expected strain from a waveform model. The vector $\bm{\theta}_{\rm GW}$ includes parameters that are particularly useful to infer EoS properties, $\bm{\theta}^{\rm EoS}_{\rm GW}$, and nuisance parameters, $\bm{\theta}^{\rm nuis.}_{\rm GW}$, that are essential when performing analysis on GW-emitting binaries. Including a dozen nuisance parameters, however, significantly slows down the sampling process and hence we calculate the likelihood through the high-precision interpolation in TOAST [112] that marginalizes over these parameters,
@@
where $\mathcal{M}=(M_{1}M_{2})^{3/5}/(M_{1}+M_{2})^{1/5}$ is the chirp mass, $q=M_{1}/M_{2}$ is the mass ratio of the event $i$, and $\Lambda_{1}(M_{1})$ and $\Lambda_{2}(M_{2})$ the TDs of the individual stars. The TD is related to the component star’s mass and radius through the underlying EoS and its radial profile.
@@ III.2.4 HESS J1731-347
In the case of unusually light neutron star HESS J1731-347 (hereafter J1731), a consensus on its gross parameters has not been reached yet on the basis of observational modeling. Nevertheless, we tentatively incorporate this source into our analysis for theoretical exploration. Specifically, we adopt one of the two most reliable results from Ref. [64], which utilized the most advanced physical modeling and the most precise distance measurements available. Similar to the NICER sources, we apply the KDE method to the released model samples to generate the posterior distributions, which are then treated as the likelihood $\mathcal{L}_{\rm{HESS}}(\bm{\theta}_{\rm{EoS}})$.
Figure 5: Posteriors for mass-radius (upper panels) and mass-tidal deformability (lower panels) distributions under different astrophysical scenarios explored in this work. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B (left panels) and F (right panels). The various color solid lines show the results at 95.4% CI for scenarios which take additionally into account the mass-radius estimates of HESS J1731 (B2 and F2), the mass measurement of PSR J0952 (B3 and F3), or both of these masurements (B4 and F4). Finally, the dash-dotted lines correspond to the scenarios which do not include the $\chi$EFT constraints. In the lower panels, the insets magnify the TDs for typical masses of CSs involved in GW events. In the upper panels, the elliptical contours of the input mass-radius estimates of HESS J1731 and the band for the mass of PSR J0952 are shown as well for reference.
Figure 6: Probability distribution functions (PDFs) for maximum mass $M_{\rm max}$ under different astrophysical scenarios. The shaded regions represent the distributions for scenarios B1 (left panel) and F1 (right panel), while the various color solid lines correspond to the scenarios which take additionally into account the mass-radius estimates of HESS J1731 (B2 and F2), mass measurement of PSR J0952 (B3 and F3), or both of these two measurements (B4 and F4). The dash-dotted lines correspond to neglecting the $\chi$EFT constraints in scenarios B1 and F1.
Figure 7: Correlation between the radii $R_{2.0}$ and $R_{1.4}$ (upper panels), and the maximum mass $M_{\rm max}$ and the radius-difference $\Delta R=R_{1.4}-R_{2.0}$ (lower panels) under different astrophysical scenarios. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B1 (left panels) and F1 (right panels), while the various color solid lines correspond to 95.4% CI for the scenarios which take additionally into account the mass-radius estimates of HESS J1731 (B2 and F2), mass measurement of PSR J0952 (B3 and F3), or both of these two measurements (B4 and F4). The dash-dotted line for scenarios (B1 and F1) correspond to neglecting the $\chi$EFT constraints.
@@ IV Multimessenger inferences on compact stars and dense matter
We apply the theoretical and astrophysical constraints both selectively and in concert to elucidate the importance of the individual impacts of the ultra-compact low-mass star (HESS J1731) and the very high mass star (PSR J0952) as well as the overall picture that emerges from a combination of such constraints.
Before embarking onto detailed discussion we present the scenarios that we consider in Table 4. Specifically, we will examine two baseline scenarios, labeled B1 and F1, which share the same data for PSR J0437 and PSR J0740 but incorporate two different analyses for PSR J0030 using different models of the surface temperature patterns, and two different analyses for PSR J1231 using different radius priors. For each baseline scenario, we implement further the data for the ultra-compact low-mass star HESS J1731 (denoted as scenarios B2 and F2), and for the high-mass star PSR J0952 (denoted as scenarios B3 and F3). Finally we take into account both constraints (denoted as scenarios B4 and F4) in our analysis.
@@ IV.1 Implications for properties of compact stars
Figure 8: Posteriors for the nucleonic compact-star EoS (upper panels) distribution and the proton fraction (lower panels) under different scenarios. The shaded regions represent the distributions at 68.3% and 95.4% CIs for scenarios B1 (left panels) and F1 (right panels), while the various color solid lines correspond to 95.4% CI for the scenarios which take additionally into account the mass-radius estimates of HESS J1731 (B2 and F2), the mass measurement of PSR J0952 (B3 and F3), and both of these two measurements (B4 and F4). The dash-dotted line for scenarios (B1 and F1) corresponds to neglecting the $\chi$EFT constraint. In each panel, the contours show the corresponding distributions of the respective $2.0\,M_{\odot}$ and the maximum-mass configurations for scenarios B1 (B4) and F1 (F4). In lower panels the orange bands labeled $F_{\rm DU}$ show the admissible threshold values for the onset of direct Urca (DU) cooling process.
@@
Incorporating the mass and radius estimates for HESS J1731 results in only a slight shift—approximately 0.1 km—to the lower-radius side in the posterior region for scenario B2 (F2) compared to B1 (F1); see Fig. 5. As expected, the posterior distributions for tidal deformability ($\Lambda$) exhibit similar trends to the $M$-$R$ posteriors due to the strong positive correlation between $\Lambda$ and $R$ for a given mass (e.g., Ref [47] and references therein). Consequently, the posterior region overlaps with the mass and radius estimates for HESS J1731 only at the 95.4% CI. Meanwhile, the impact on the maximum mass ($M_{\rm max}$) distribution remains negligible in scenarios B1 and B2 (F1 and F2); see Fig. 6. This suggests that the mass and radius measurements of four pulsars from NICER, combined with the TDs inferred from GW170817, GW190425 and the constraints from $\chi$EFT calculations, primarily define the posterior region of the $M$-$R$ distribution. The effect of the $\chi$EFT constraint on the $M$-$R$ posterior is particularly evident in scenarios B1 and F1 (Fig. 5), where it imposes a lower bound on the radii of sub-canonical-mass stars. Additionally, the mass and radius inferences derived for HESS J1731 shows only weak overlap with the $M$-$R$ ellipse for PSR J1231 (Fig. 1), which limits the influence of HESS J1731 data in distinguishing viable EoSs.
@@
Including the mass measurement for PSR J0952 instead, namely for scenario B3 (F3), the posterior region is shifted to a larger radius region by 0.1-0.4 km compared to scenario B1 (F1). A more significant difference appears for stars with mass $M>1.4\,M_{\odot}$. This is because the radii for sub-canonical-mass stars are well-constrained, while the NICER inferences for PSR J0030 and J0740, allow for radii larger than $14$ km, which then allow stiffer high-density EoSs. As anticipated, the maximum mass $M_{\rm max}$ is enlarged by $0.1\,M_{\odot}$, with, in particular, a higher probability density for scenario F3; see Fig. 6. The maximum mass now is $M_{\rm max}=2.28_{-0.19}^{+0.20}\,M_{\odot}$ and $2.39_{-0.23}^{+0.19}\,M_{\odot}$ respectively for scenarios B3 and F3. Meanwhile, the distributions of maximum mass $M_{\rm max}$ is somewhat narrowed compared to scenarios B1 and F1, respectively.
@@
Finally, incorporating both HESS J1731 and PSR J0952 data, namely for scenarios B4 and F4, the boundaries for the posterior regions are located well within the ranges set by the two cases discussed above. The distributions of maximum mass $M_{\rm max}$ are rather close to those of scenarios B3 and F3, respectively. To conclude, the $M$-$R$ estimates for HESS J1731 and mass measurement for PSR J0952 affect only the finer details of the posterior compared to the case where these constraints are not included.
@@
Reference [113] pointed out that the sign of the difference of the radii of CSs of masses 1.4 and $2.0\,M_{\odot}$, $\Delta R=R_{1.4}-R_{2.0}$ is an indicator that the EoS softens (if positive) or stiffens (if negative) at high densities. Figure 7 illustrates the posterior distributions for the correlation between $R_{1.4}$ and $R_{2.0}$, and the correlation between the maximum mass $M_{\rm max}$ and the radius difference $\Delta R$. It is seen that the correlations between these parameters are similar across all scenarios. For scenarios B (B refers to B1–B4), the preference is for $\Delta R$ being positive, i.e., $R_{1.4}>R_{2.0}$, but for scenarios F, negative $\Delta R$ is allowed even at 68.3% CI. More details of correlations among properties of CSs will be discussed below in Subsec. IV.3. The predicted key quantities of CSs under different astrophysical scenarios are summarized in the Appendix.